OpenStudy (anonymous):

Given f(x)= x -10. What is f(-7)? A. -17 B. -10 C. -7 D. -3 Please help me with the answer ill give u fan + medal!

OpenStudy (amorfide):

you have a function f in terms of x f(x)=x-10 f(-7) is a function where x is equal to -7 so substitute x=-7

rishavraj (rishavraj):

just plug x = -7 in the equation...and get the answer

OpenStudy (anonymous):

whats my answer i dont get it

rishavraj (rishavraj):

see the equation is x - 10 plug x = -7 in tht equation.....

sammixboo (sammixboo):

Well if x = -7 then plug it into the equation f(x) = x - 10 f(-7) = (-7) = 10 You answer is -7 - 10, which is?

OpenStudy (anonymous):

-17 is my answer?